OpenStudy (princesssleelee):

Using the completing-the-square method, find the vertex of the function f(x) = 2x2 − 8x + 6 and indicate whether it is a minimum or a maximum and at what point. Maximum at (2, –2) Minimum at (2, –2) Maximum at (2, 6) Minimum at (2, 6)

OpenStudy (princesssleelee):

im stuck between a and be, i know those points are correct, however i do not know how to determine max. & min.

OpenStudy (solomonzelman):

Can you show the work that you have so far?

OpenStudy (triciaal):

show your work to complete the square

OpenStudy (princesssleelee):

its all on my scrap paper.

OpenStudy (princesssleelee):

believe me, you wouldnt be able to understand it lol

OpenStudy (princesssleelee):

i cant even read it...

OpenStudy (solomonzelman):

Can you type (at least some of) it up here?

OpenStudy (triciaal):

so just do it again here

OpenStudy (princesssleelee):

i dont know what i did honestly my nephew helped with mostly all of it

OpenStudy (solomonzelman):

Are you aware of the property below? \(\color{#000000 }{ \displaystyle x^2+2ax+a^2=(x+a)^2 }\)

OpenStudy (princesssleelee):

barely.

OpenStudy (solomonzelman):

Please expand the expression below: \(\color{#000000 }{ \displaystyle (x+a)\cdot (x+a) }\)

OpenStudy (princesssleelee):

i got it! thank you very much! i came up with answer choice B. he came back with the paperwork and explained to me what i jotted down so quickly lol.

OpenStudy (solomonzelman):

Yes, B is right!

OpenStudy (princesssleelee):

Thank you so much!

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(x)=2x^2-8x+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2-4x\right)+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2-4x+0\right)+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2-4x\color{blue}{+4}\color{red}{-4}\right)+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2-4x\color{blue}{+4}\right)+2\cdot(\color{red}{-4})+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x^2-2\cdot(2) x+2^2\right)-8+6}\) \(\color{#000000 }{ \displaystyle f(x)=2\left(x-2\right)^2-2}\) \(\color{#000000 }{ \displaystyle (h,k)=(2,-2)}\) is the vertex. (And since parabola opens up, the vertex is minimum)